OpenStudy (piercetheveil47):

Choose the equation below that represents the line passing through the point (1, -4) with a slope of one half.

OpenStudy (piercetheveil47):

y - 4 = one half(x - 1) y + 4 = 2(x - 1) y + 4 = one half(x - 1) y - 4 = 2(x + 1)

OpenStudy (piercetheveil47):

heeelp :(

OpenStudy (adamaero):

what are your thoughts?

OpenStudy (adamaero):

http://www.purplemath.com/modules/strtlneq.htm

OpenStudy (piercetheveil47):

i think it may be A or C?

OpenStudy (adamaero):

yes, but why?

OpenStudy (piercetheveil47):

I think it is A or C bc you need to multiply the 1/2

OpenStudy (adamaero):

so you know that the "m" in this equation represents the slope: y = mx + b

OpenStudy (piercetheveil47):

yes

OpenStudy (piercetheveil47):

i just don't know how to figure out if they are adding or subtracting 4

OpenStudy (adamaero):

If you still don't know the answer from that picture as a guide, say so

OpenStudy (piercetheveil47):

so that means it is C?

OpenStudy (adamaero):

why C?

OpenStudy (piercetheveil47):

because it says that it is negative 4?

OpenStudy (adamaero):

well, ok--close enough--you're correct

OpenStudy (piercetheveil47):

okay thank you! @adamaero